OpenStudy (anonymous):

Add four and one-sixth plus three and one-third. seven and one-half......i think this one seven and two-thirds eight and one-half eight and two-thirds

OpenStudy (anonymous):

seven and one-half......i think this one

OpenStudy (anonymous):

4 1/6 + 3 1/3 = 4 1/6 + 3 2/6 = 7 3/6 = 7 1/2